6. The slant height and base diameter of a conical tomb are
25 m and 14 m respectively. Find the cost of white­
washing iJs curved surface at the rate of r 210 per 100 m2

Grade:12th pass

2 Answers

Pawan Prajapati
3 years ago
. Curved surface area = rcrl = 22 x 7 x 26 m2 = 660 m2 7 Cost of white washing = f (210 x ) = f 1165.
Ram Kushwah
3 years ago
For the conical tomb given that :
l=25 m ,diameter d=14 m
so r=14/2=7 m
We know that the the curved surface area of a cone=πrl
=22/7 x 7 x 25
=25 x 22
=550 m²
Now rate of painting= Rs 210 /100 m²
Hence the painting charges are:
=210x550/100
=21x55
=Rs 1155
Thus the pianting charge of the tomb will be Rs 1155